OpenStudy (anonymous):

Which ratio can be used to form a proportion with the given ratio? 8 to 12 A. 6 to 9 B. 10 to 14 C. 12 to 16 D. 64 to 144

ganeshie8 (ganeshie8):

u need to test each option

OpenStudy (anonymous):

help me

ganeshie8 (ganeshie8):

lets test and see if first options forms the proportion :- \(\large \frac{8}{12} = \frac{6}{9}\)

ganeshie8 (ganeshie8):

cross multiply, and see if u get samething both

ganeshie8 (ganeshie8):

cross multiply, and see if u get samething both sides

OpenStudy (anonymous):

i did not get it

OpenStudy (shamil98):

You would reduce the fraction, or just cross multiply to see if the result is the same on both sides.

OpenStudy (anonymous):

can u show me

OpenStudy (shamil98):

\[\text{Does}~~\frac{ 8 }{ 12 } = \frac{ 6 }{ 9 }~ ?\] Reduce the fractions you get.. \[\frac{ 2 }{ 3 } = \frac{ 2 }{ 3 }\]

OpenStudy (anonymous):

k so